# Stability of Non-Hyperbolic Sets

### v1

@@ 8 Finite Collet-Eckmann Condition
In this section we set up the induction step needed to prove later the existence of maps with a Collet-Eckmann point. More in details we prove the existence of maps having a point satisfying the finite Collet-Eckmann condition. Roughly speaking a point $x$ is a $T$-Collet-Eckmann (T-CE) point if there are vectors $v$ in the tangent space at $x$ such that the derivative of the map along the orbit of $x$ in the direction of $v$ grows at the least exponentially fast up to time $T$. Below is the formal definition.
Definition 11.
Let $f:\mathcal{M}\to\mathcal{M}$ be a map and let $x\in\mathcal{M}$. Given constants $\varphi>0$, $0<C<1$, $\rho>1$, we say that $x$ is a $T$-Collet-Eckmann (T-CE) point if the following holds. Denote by $K_{\varphi}(x)$ the cone at $x$ of angle $\varphi$}, and by $\text{T}_{x}$ the tangent space at $x$, then for all vectors $v\in K_{\varphi}\subset\text{T}_{x}$,
@@
$\lambda^{\theta}\mu^{1+\alpha}>\lambda^{\alpha},$ (8.1)
where $\alpha$ is defined in (4.17). By (2.17) and (4.17), there exists $\theta\in(\theta_{0},\theta_{1})$ satisfying (8.1), see Appendix.
Take constants $\varphi>0$, $0<C<1$ and $\rho>1$ such that,
@@
We introduce now some notation. Let $x$ be a point. If $\Gamma$ is a curve, we denote by $\text{d}_{\text{v}}\left(x,\Gamma\right)$ the vertical distance between a point $x$ and the curve $\Gamma$. The angle between two vectors is denoted by $\angle$.
Let $S$ be a tangency strip and let $\left\{b,U\right\}$ be a $(1,N,M)$ descendant of $S$. For all $f\in U$, let $\mathfrak{D}(f)$ be the $\left(1,N,M\right)$ dynamical partition of $f$. Given $T\in\left\{1,\dots,N-1\right\}$, one can define a vector field $\mathfrak{X}\left(f^{T}\right)$ on $C^{N-T}$ as
@@
Proposition 8.
Let $S$ be a tangency strip and let $\left\{b,U\right\}$ be a $(1,N,M)$ descendant of $S$. Suppose that for all $f\in U$, there exist a positive constant $K$ and $T\in\left\{1,\dots,N-1\right\}$ such that the following holds,
@@
      $Df_{x}^{N-T}e_{2}\in\text{span}\left(\mathfrak{X}\left(f^{N-T}\right)\right)\iff x\in\Gamma.$
@@
      $\frac{1}{K}{{d}_{\text{v}}}\left(x,\Gamma\right)\leq\angle\left(Df_{x}^{N-T}e_{2},\mathfrak{X}\left(f^{N-T}\right)\right)\leq Kd_{\text{v}}\left(x,\Gamma\right).$
@@
      $Df^{\tilde{N}-\tilde{T}}(x)e_{2}\in\text{span}\left({\mathfrak{X}}\left(f^{\tilde{N}-\tilde{T}}\right)\right)\iff x\in\tilde{\Gamma}$.
@@
[…]
$\tilde{T}=T+n+N-T+T+n_{0}+M+\overline{\ell},$
and for all $\tilde{f}\in\tilde{U}$ let $\tilde{K}$ be a positive constant assuring that $\text{IH}_{\tilde{H}^{0}}$, $\text{IH}_{\tilde{C}^{0}}$-(iii) and $\text{IH}_{\tilde{C}^{\tilde{N}-\tilde{T}}}-$(iii) hold. The existence of such a constant is guaranteed by the initial conditions on the family. Observe that $\text{IH}_{\tilde{Q}}$ is trivially satisfies.
Define $\tilde{\Gamma}=f^{-\overline{\ell}}\left(\Gamma\right)$, then $\text{IH}_{\tilde{C}^{0}}$-(i) and $\text{IH}_{\tilde{C}^{0}}$-(ii) are satisfied by construction.
The rest of the proof is devoted to prove $\text{IH}_{\tilde{C}^{\tilde{N}-\tilde{T}}}-$(i) and $\text{IH}_{\tilde{C}^{\tilde{N}-\tilde{T}}}-$(ii). This will be achieved by steps. Each step corresponds to a time defining $\tilde{T}$. In order to lighten the notation, we denote $\tilde{C}=\tilde{C}^{\tilde{N}-\tilde{T}}$. Let $x\in\tilde{C}$ and let $v\in K_{\varphi}(x)\subset\text{T}_{x}$. Observe that $\tilde{C}\subset C^{N-T}$.
@@
    We begin by proving $\text{IH}_{\tilde{C}^{\tilde{N}-\tilde{T}}}-$(i) from time 0 up to time $T$. By $\text{IH}_{C^{N-T}}$-(i),
    • $\text{IH}_{\tilde{C}}-$(i)P1:
      $\left|Df_{x}^{s}v\right|\geq C\rho^{s}\left|v\right|\text{ }\forall\text{ }0\leq s\leq T.$
@@
    $v^{(1)}_{2}>\cos\varphi\left|v^{(1)}\right|\text{ and }\frac{v^{(1)}_{1}}{v^{(1)}_{2}}<\tan\varphi.$ (8.3)
@@
    We begin by proving $\text{IH}_{\tilde{C}^{\tilde{N}-\tilde{T}}}-$(i) from time $T+1$ up to time $T+n$. Indeed, by $\text{IH}_{Q}$ and $\text{IH}_{\tilde{C}}-$(i)P1, for all $1\leq k\leq n$, we get,
    • $\text{IH}_{\tilde{C}}-$(i)P2:
      $\left|Df^{T+k}_{x}v\right|$ $=$ $\left|\left(\begin{matrix}\lambda^{k}&0\\ 0&\mu^{k}\end{matrix}\right)\left(\begin{matrix}v^{(1)}_{1}\\ v^{(1)}_{2}\end{matrix}\right)\right|>\mu^{k}v^{(1)}_{2}\geq\mu^{k}\cos\varphi\left|v^{(1)}\right|$
      $=$ $\mu^{k}\cos\varphi\left|Df^{T}_{x}v\right|\geq\mu^{k}\cos\varphi C\rho^{T}\left|v\right|\geq C\rho^{T+k}\left|v\right|.$
@@
    $\mu^{k}\cos\varphi\geq\rho^{k},\text{ }\forall 1\leq k\leq n.$
@@
    $\left|Df^{T+n}_{x}v\right|\geq\mu^{n}\cos\varphi C\rho^{T}\left|v\right|.$ (8.4)
@@
    $v^{(2)}=Df^{T+n}_{x}v=\left(\begin{matrix}v^{(2)}_{1}\\ v^{(2)}_{2}\end{matrix}\right)$.
    By $\text{IH}_{Q}$ and by (8.3),
    $\frac{v^{(2)}_{1}}{v^{(2)}_{2}}=\left(\frac{\lambda}{\mu}\right)^{n}\frac{v^{(1)}_{1}}{v^{(1)}_{2}}<\left(\frac{\lambda}{\mu}\right)^{n}\tan\varphi=\left(\frac{\lambda}{\mu}\right)^{n}\varphi.$ (8.5)
@@
    We start proving $\text{IH}_{\tilde{C}^{\tilde{N}-\tilde{T}}}-$(i) from time $T+n+1$ up to time T+n+NTT+n+N-T. Indeed, by $\text{IH}_{C^{0}}$-(iii) and (8.4), for all $1\leq k\leq N-T$, we get,
    • $\text{IH}_{\tilde{C}}-$(i)P3:
      $\left|Df^{T+n+k}_{x}v\right|\geq\frac{1}{K}\left|Df^{T+n}_{x}v\right|>\frac{1}{K}\mu^{n}\cos\varphi C\rho^{T}\left|v\right|\geq C\rho^{T+n+k}\left|v\right|.$
@@
    $\frac{1}{K}\left(\frac{\mu}{\rho}\right)^{n}\cos\varphi\geq\rho^{k},\text{ }\forall 1\leq k\leq N-T$
@@
    $\frac{1}{K}\left(\frac{\mu}{\rho}\right)^{n}\cos\varphi\geq\rho^{N-T}.$
    This follow from the condition imposed in (8)-(oo) and taking $n$ large enough. In particular, $\text{IH}_{\tilde{C}}-$(i)P3 proves
    $\left|Df^{T+n+N-T}_{x}v\right|\geq\frac{C}{K}cos\varphi\mu^{n}\ \rho^{T}\left|v\right|.$ (8.6)
    Moreover by $\text{IH}_{C^{0}}$-(ii) and Lemma 7,
    $\angle\left(Df_{f^{T+n}(x)}^{N-T}e_{2},\mathfrak{X}\left(f^{T+n+N-T}\right)\right)\geq\frac{1}{K}{\text{d}_{\text{v}}}\left(f^{T+n}(x),\Gamma\right)\geq\frac{1}{K^{2}}\left(\lambda^{\theta}\mu\right)^{n},$
    and by (8.5),
@@
    The last two inequalities together with $\text{IH}_{C^{0}}$-(iii), gives that, for $n$ large enough,
    $\angle\left(Df_{x}^{T+n+N-T}v,\mathfrak{X}\left(f^{T+n+N-T}\right)\right)\geq\frac{1}{2K^{2}}\left(\lambda^{\theta}\mu\right)^{n}.$ (8.7)
@@
    We start proving $\text{IH}_{\tilde{C}^{\tilde{N}-\tilde{T}}}-$(i) from time T+nNT+1T+nN-T+1 up to time T+n+NT+TT+n+N-T+T. Indeed, by $\text{IH}_{C^{N-T}}$-(iii) and (8.6), for all $1\leq k\leq T$, we get,
    • $\text{IH}_{\tilde{C}}-$(i)P4:
      $\left|Df^{T+n+N-T+k}_{x}v\right|\geq\frac{1}{K}\left|Df^{T+n+N-T}_{x}v\right|\geq\frac{C}{K^{2}}\cos\varphi\mu^{n}\rho^{T}\left|v\right|\geq C\rho^{T+n+N-T+k}\left|v\right|.$
@@
    $\frac{1}{K^{2}}\cos\varphi\left(\frac{\mu}{\rho}\right)^{n}\rho^{T}\geq\rho^{N+k},\text{ }\forall 1\leq k\leq T$
@@
    $\frac{1}{K^{2}}\cos\varphi\left(\frac{\mu}{\rho}\right)^{n}\geq\rho^{N}.$
    This follow from the condition imposed in (8)-(oo) and taking $n$ large enough. In particular, $\text{IH}_{\tilde{C}}-$(i)P4 proves
    $\left|Df^{T+n+N}_{x}v\right|\geq\frac{C}{K^{2}}cos\varphi\mu^{n}\ \rho^{T}\left|v\right|.$ (8.8)
    Equation (8.7), $\text{IH}_{C}$-(iii) and the definition of the vector field $\mathfrak{X}$ give that, for $n$ large enough,
    $\angle\left(Df_{x}^{T+n+N}v,e_{1}\right)\geq\frac{1}{2K^{3}}\left(\lambda^{\theta}\mu\right)^{n}.$ (8.9)
@@
    $v^{(4)}=Df^{T+n+N}_{x}v=\left(\begin{matrix}v^{(4)}_{1}\\ v^{(4)}_{2}\end{matrix}\right)$,
    then, by (8.9),
    $\frac{v^{(4)}_{2}}{\left|v^{(4)}\right|}\geq\sin\frac{1}{2K^{3}}\left(\lambda^{\theta}\mu\right)^{n}\geq\frac{1}{4K^{3}}\left(\lambda^{\theta}\mu\right)^{n}$ (8.10)
@@
    We start proving $\text{IH}_{\tilde{C}^{\tilde{N}-\tilde{T}}}-$(i) from time $T+nN+1$ up to time $T+n+N+n_{0}$. Indeed, by $\text{IH}_{Q}$, (8.10) and (8.8), for all $1\leq k\leq n_{0}$, we get,
    • $\text{IH}_{\tilde{C}}-$(i)P5:
      $\left|Df^{T+n+N+k}_{x}v\right|$ $\geq$ $\mu^{k}v^{(4)}_{2}\geq\mu^{k}\frac{1}{4K^{3}}\left(\lambda^{\theta}\mu\right)^{n}\left|Df^{T+n+N}_{x}v\right|$
      $\geq$ $\frac{C}{4K^{5}}\cos\varphi\mu^{k+n}\rho^{T}\left(\lambda^{\theta}\mu\right)^{n}\left|v\right|\geq C\rho^{T+n+N+k}\left|v\right|.$
@@
    $\frac{1}{4K^{5}}\cos\varphi\left(\frac{\mu}{\rho}\right)^{n+k}\left(\lambda^{\theta}\mu\right)^{n}\geq\rho^{N},\text{ }\forall 1\leq k\leq n_{0}$
@@
    $\frac{1}{4K^{5}}\cos\varphi\left(\frac{\mu}{\rho}\right)\left(\lambda^{\theta}\mu^{2}\right)^{n}\geq\rho^{n+N}.$
    This follow from the condition imposed in (8)-(oo) and taking $n$ large enough. In particular, $\text{IH}_{\tilde{C}}-$(i)P5 proves
    $\left|Df^{T+n+N+n_{0}}_{x}v\right|\geq\frac{C}{4K^{5}}\cos\varphi\mu^{n_{0}}\rho^{T}\left(\lambda^{\theta}\mu^{2}\right)^{n}\left|v\right|.$ (8.11)
    Moreover, by (8.9),
    $\tan\angle\left(Df_{x}^{T+n+N}v,e_{2}\right)\leq\frac{2K^{3}}{\left(\lambda^{\theta}\mu\right)^{n}}$
    which together with (8.1) and by taking $n$ large enough imply that,
    $\tan\angle\left(Df_{x}^{T+n+N+n_{0}}v,e_{2}\right)\leq\left(\frac{\lambda}{\mu}\right)^{n_{0}}\frac{2K^{3}}{\left(\lambda^{\theta}\mu\right)^{n}}<<1.$ (8.12)
@@
    We start proving $\text{IH}_{\tilde{C}^{\tilde{N}-\tilde{T}}}-$(i) from time $T+nN+n_{0}+1$ up to time $T+n+N+n_{0}+M$. Indeed, by $\text{IH}_{H}$ and (8.11), for all $1\leq k\leq M$, we get,
    • $\text{IH}_{\tilde{C}}-$(i)P6:
      $\left|Df^{T+n+N+n_{0}+k}_{x}v\right|$ $\geq$ $\frac{1}{K}\left|Df^{T+n+N+n_{0}}_{x}v\right|$
      $\geq$ $\frac{C}{4K^{6}}\cos\varphi\mu^{n_{0}}\rho^{T}\left(\lambda^{\theta}\mu^{2}\right)^{n}\left|v\right|$
      $\geq$ $C\rho^{T+n+N+n_{0}+k}\left|v\right|.$

### v3

@@ 8 Finite Collet-Eckmann Condition
In this section we describe the induction step needed to prove later the existence of maps with a Collet-Eckmann point. More in detail, we prove the existence of maps having a point satisfying the finite Collet-Eckmann condition. Roughly speaking a point $x$ is a $T$-Collet-Eckmann (T-CE) point if there are vectors $v$ in the tangent space at $x$ such that the derivative of the map along the orbit of $x$ in the direction of $v$ grows at the least exponentially fast up to time $T$. Below is the formal definition.
As before, without lose of generality, we consider a diffeomorphism $f:\mathcal{M}\to\mathcal{M}$ where the manifold $\mathcal{M}$ is an open set in $\mathbb{R}^{2}$. Let $\varphi>0$. Denote by $\text{T}_{x}\mathcal{M}$ the tangent space at $x$, and by
$K_{\varphi}(x)=\left\{(v_{1},v_{2})\in\text{T}_{x}\mathcal{M}\left|\right.|v_{2}|\tan\varphi>|v_{1}|\right\},$
the cone at $x$ of angle $\varphi$}.
Definition 8.0.1.
Given constants $0<C<1$, $\rho>1$, we say that $x\in\mathcal{M}$ is a $T$-Collet-Eckmann (T-CE) point if the following holds. For all vectors $v\in K_{\varphi}(x)$,
@@
$\lambda^{\theta}\mu^{1+\alpha}>\lambda^{\alpha},$ (8.0.2)
where $\alpha$ is defined in (4.1.49). By (2.3.2) and (4.1.49), there exists $\theta\in(\theta_{0},\theta_{1})$ satisfying (8.0.2), see Appendix. Take constants $\varphi>0$, $0<C<1$ and $\rho>1$ such that,
@@
We introduce now some notation. Let $(x,y)$ be a point and let $\Gamma$ be a curve which is a graph over the $x$-axes. If there is a point $(x,y_{0})\in\Gamma$ then denote by $\text{d}_{\text{v}}\left(x,\Gamma\right)=|y-y_{0}|$ the vertical distance between a point $x$ and the curve $\Gamma$. The angle between two vectors is denoted by $\angle$.
Let $S$ be a tangency strip and let $\left\{b,U\right\}$ be a $(1,N,M)$ descendant of $S$. For all $f\in U$, let $\mathfrak{D}(f)$ be the $\left(1,N,M\right)$ dynamical partition of $f$. Given $T\in\left\{1,\dots,N-1\right\}$, one can define a vector field $\mathfrak{X}^{N-T}$ on $C^{N-T}$ as
@@
Proposition 8.0.4.
Let $S$ be a $r$-tangency strip and let $\left\{b,U\right\}$ be a $(1,N,M)$ descendant of $S$. Suppose that for all $f\in U$, there exist a positive constant $K$ and $T\in\left\{1,\dots,N-1\right\}$ such that the $T$-Collet-Eckmann Condition holds,
@@
      $Df_{x}^{N-T}e_{2}\in\text{span}\left(\mathfrak{X}^{N-T}\right)\iff x\in\Gamma.$
@@
      $\frac{1}{K}{{d}_{\text{v}}}\left(x,\Gamma\right)\leq\angle\left(Df_{x}^{N-T}e_{2},\mathfrak{X}^{N-T}\right)\leq Kd_{\text{v}}\left(x,\Gamma\right).$
@@
$\tilde{C}^{\tilde{N}-\tilde{T}}\subset C^{N-T}$
such that the $\tilde{T}-$Collet-Eckmann Condition holds. Namely,
@@
      $Df^{\tilde{N}-\tilde{T}}(x)e_{2}\in\text{span}\left({\mathfrak{X}}^{\tilde{N}-\tilde{T}}\right)\iff x\in\tilde{\Gamma}$.
@@
Let $\tilde{S}=S_{j}\subset U\subset S$ be a tangency strip and let $\left\{\tilde{b},\tilde{U}\right\}=\left\{b_{k}^{(j)},U_{k}^{(j)}\right\}$ be a $\left(1,\tilde{N},\tilde{M}\right)=\left(1,N_{k}^{(j)},M_{k}^{(j)}\right)$ descendant of $\tilde{S}$ as in Proposition 7.4.1. In each step of the construction we might have to increase the $j$. Consider the orbit from $\tilde{C}^{0}$ to $\tilde{C}^{\tilde{N}}$. By construction there exist exactly three moments, $s_{1}$, $s_{2}$, $s_{3}$ smaller than $\tilde{N}$ such that $\tilde{C}^{s_{i}}\subset C^{N-T}$. Using the notation from Remark 7.4.4 we get
$s_{1}$ $=$ $\overline{\ell}_{k}^{(j)}+N-T,$
$s_{2}$ $=$ $s_{1}+T+\theta n_{k}^{(j)}+N-T,$
$s_{3}$ $=$ $s_{2}+T+n_{k}^{(j)}+N-T.$
Define $\tilde{T}$ such that $\tilde{N}-\tilde{T}=s_{2}$ corresponds to the second passage [^1] The second passage is an essential choice. Using the third passage will not lead to a Cantor set, see Lemma 9.3.5 and its proof. Using the first passage will not give enough expansion to ensure the CE condition. trough $C^{N-T}$, see Figure 12.
Figure 12: Passages in $C^{N-T}$
In particular,
$\tilde{T}=T+n_{k}^{(j)}+N-T+T+n_{0,k}^{(j)}+M+\underline{\ell}_{k}^{(j)}.$
For all $\tilde{f}\in\tilde{U}$, let $\tilde{K}$ be a positive constant describing the bounds as required in $\text{IH}_{\tilde{H}^{0}}$, $\text{IH}_{\tilde{C}^{0}}$-(iii) and $\text{IH}_{\tilde{C}^{\tilde{N}-\tilde{T}}}-$(iii). The existence of such a constant is guaranteed by the fact that this conditions consider only a finite number of iterates of the original diffeomorphism. Observe that $\text{IH}_{\tilde{Q}}$ is trivially satisfies. Let $\tilde{\Gamma}$, be the curve associated to the critical point $\tilde{c}$, then $\text{IH}_{\tilde{C}^{0}}$-(i) and $\text{IH}_{\tilde{C}^{0}}$-(ii) are satisfied by construction.
The rest of the proof is devoted to prove $\text{IH}_{\tilde{C}^{\tilde{N}-\tilde{T}}}-$(i) and $\text{IH}_{\tilde{C}^{\tilde{N}-\tilde{T}}}-$(ii). This will be achieved by steps. Each step corresponds to a term in the definition of $\tilde{T}$. In order to lighten the notation, we denote $\tilde{C}=\tilde{C}^{\tilde{N}-\tilde{T}}$. Let $x\in\tilde{C}$ and let $v\in K_{\varphi}(x)\subset\text{T}_{x}$. Observe that $\tilde{C}\subset C^{N-T}$.
@@
    We begin by proving $\text{IH}_{\tilde{C}^{\tilde{N}-\tilde{T}}}-$(i) from time 0 up to time $T$. By $\text{IH}_{C^{N-T}}$-(i), we get,
    $\text{IH}_{\tilde{C}}\text{-(i)P1: }\left|Df_{x}^{s}v\right|\geq C\rho^{s}\left|v\right|\text{ }\forall\text{ }0\leq s\leq T.$
@@
    $v^{(1)}_{2}>\cos\varphi\left|v^{(1)}\right|\text{ and }\frac{v^{(1)}_{1}}{v^{(1)}_{2}}<\tan\varphi.$ (8.0.5)
@@
    Now we prove $\text{IH}_{\tilde{C}^{\tilde{N}-\tilde{T}}}-$(i) from time $T+1$ up to time $T+n_{k}^{(j)}$. Indeed, by $\text{IH}_{Q}$, $\text{IH}_{\tilde{C}}-$(i)P1 and (8.0.5), for all $1\leq s\leq n$, we get,
    $\text{IH}_{\tilde{C}}\text{-(i)P2: }\left|Df^{T+s}_{x}v\right|$ $=$ $\left|\left(\begin{matrix}\lambda^{s}&0\\ 0&\mu^{s}\end{matrix}\right)\left(\begin{matrix}v^{(1)}_{1}\\ v^{(1)}_{2}\end{matrix}\right)\right|>\mu^{s}v^{(1)}_{2}\geq\mu^{s}\cos\varphi\left|v^{(1)}\right|$
    $=$ $\mu^{s}\cos\varphi\left|Df^{T}_{x}v\right|\geq\mu^{s}\cos\varphi C\rho^{T}\left|v\right|\geq C\rho^{T+s}\left|v\right|.$
@@
    $\mu^{s}\cos\varphi\geq\rho^{s},\text{ }\forall 1\leq s\leq n.$
@@
    $\left|Df^{T+n_{k}^{(j)}}_{x}v\right|\geq\mu^{n_{k}^{(j)}}\cos\varphi C\rho^{T}\left|v\right|.$ (8.0.6)
@@
    $v^{(2)}=Df^{T+n_{k}^{(j)}}_{x}v=\left(\begin{matrix}v^{(2)}_{1}\\ v^{(2)}_{2}\end{matrix}\right)$.
    By $\text{IH}_{Q}$ and by (8.0.5),
    $\frac{v^{(2)}_{1}}{v^{(2)}_{2}}=\left(\frac{\lambda}{\mu}\right)^{n}\frac{v^{(1)}_{1}}{v^{(1)}_{2}}<\left(\frac{\lambda}{\mu}\right)^{n}\tan\varphi.$ (8.0.7)
@@
    Now we prove $\text{IH}_{\tilde{C}^{\tilde{N}-\tilde{T}}}-$(i) from time $T+n_{k}^{(j)}+1$ up to time $T+n_{k}^{(j)}+N-T$. Indeed, by $\text{IH}_{C^{0}}$-(iii) and (8.0.6), for all $1\leq s\leq N-T$, we get,
    $\text{IH}_{\tilde{C}}\text{-(i)P3: }\left|Df^{T+n_{k}^{(j)}+s}_{x}v\right|\geq\frac{1}{K}\left|Df^{T+n_{k}^{(j)}}_{x}v\right|>\frac{1}{K}\mu^{n_{k}^{(j)}}\cos\varphi\cdot C\rho^{T}\left|v\right|\geq C\rho^{T+n_{k}^{(j)}+s}\left|v\right|.$
@@
    $\frac{1}{K}\left(\frac{\mu}{\rho}\right)^{n_{k}^{(j)}}\cos\varphi\geq\rho^{s},\text{ }\forall 1\leq s\leq N-T$
@@
    $\frac{1}{K}\left(\frac{\mu}{\rho}\right)^{n_{k}^{(j)}}\cos\varphi\geq\rho^{N-T}.$
    This follow from the condition imposed in (8)-(oo) and taking $n_{k}^{(j)}$ large enough (which is equivalent to take $j$ large enough). In particular, $\text{IH}_{\tilde{C}}-$(i)P3 proves
    $\left|Df^{T+n_{k}^{(j)}+N-T}_{x}v\right|\geq\frac{C}{K}\cos\varphi\cdot\mu^{n_{k}^{(j)}}\ \rho^{T}\left|v\right|.$ (8.0.8)
    Moreover by $\text{IH}_{C^{0}}$-(ii) and Lemma 4.1.21,
    $\angle\left(Df_{f^{T+n_{k}^{(j)}}(x)}^{N-T}e_{2},\mathfrak{X}^{N-T}\right)\geq\frac{1}{K}{\text{d}_{\text{v}}}\left(f^{T+n_{k}^{(j)}}(x),\Gamma\right)\geq\frac{1}{K^{2}}\left(\lambda^{\theta}\mu\right)^{n_{k}^{(j)}},$
    and by (8.0.7),
@@
    The last two inequalities together with $\text{IH}_{C^{0}}$-(iii), gives that, for $n_{k}^{(j)}$ large enough,
    $\angle\left(Df_{x}^{T+n_{k}^{(j)}+N-T}v,\mathfrak{X}^{N-T}\right)\geq\frac{1}{2K^{2}}\left(\lambda^{\theta}\mu\right)^{n_{k}^{(j)}}.$ (8.0.9)
@@
    Now we prove $\text{IH}_{\tilde{C}^{\tilde{N}-\tilde{T}}}-$(i) from time $T+n_{k}^{(j)}+N-T+1$ up to time $T+n_{k}^{(j)}+N-T+T$. Indeed, by $\text{IH}_{C^{N-T}}$-(iii) and (8.0.8), for all $1\leq s\leq T$, we get,
    $\text{IH}_{\tilde{C}}\text{-(i)P4: }\left|Df^{T+n_{k}^{(j)}+N-T+s}_{x}v\right|$ $\geq$ $\frac{1}{K}\left|Df^{T+n_{k}^{(j)}+N-T}_{x}v\right|$
    $\geq$ $\frac{C}{K^{2}}\cos\varphi\cdot\mu^{n_{k}^{(j)}}\rho^{T}\left|v\right|\geq C\rho^{T+n_{k}^{(j)}+N-T+s}\left|v\right|.$
@@
    $\frac{1}{K^{2}}\cos\varphi\left(\frac{\mu}{\rho}\right)^{n_{k}^{(j)}}\rho^{T}\geq\rho^{N+s},\text{ }\forall 1\leq s\leq T$
@@
    $\frac{1}{K^{2}}\cos\varphi\left(\frac{\mu}{\rho}\right)^{n_{k}^{(j)}}\geq\rho^{N}.$
    This holds for $n_{k}^{(j)}$ large enough from the condition imposed in (8)-(oo). In particular, $\text{IH}_{\tilde{C}}-$(i)P4 proves
    $\left|Df^{T+n_{k}^{(j)}+N}_{x}v\right|\geq\frac{C}{K^{2}}\cos\varphi\cdot\mu^{n_{k}^{(j)}}\ \rho^{T}\left|v\right|.$ (8.0.10)
    Equation (8.0.9), $\text{IH}_{C}$-(iii) and the definition of the vector field $\mathfrak{X}$ give that, for $n_{k}^{(j)}$ large enough,
    $\angle\left(Df_{x}^{T+n_{k}^{(j)}+N}v,e_{1}\right)\geq\frac{1}{2K^{3}}\left(\lambda^{\theta}\mu\right)^{n_{k}^{(j)}}.$ (8.0.11)
@@
    $v^{(4)}=Df^{T+n_{k}^{(j)}+N}_{x}v=\left(\begin{matrix}v^{(4)}_{1}\\ v^{(4)}_{2}\end{matrix}\right)$,
    then, by (8.0.11),
    $\frac{v^{(4)}_{2}}{\left|v^{(4)}\right|}\geq\sin\frac{1}{2K^{3}}\left(\lambda^{\theta}\mu\right)^{n_{k}^{(j)}}\geq\frac{1}{4K^{3}}\left(\lambda^{\theta}\mu\right)^{n_{k}^{(j)}}$ (8.0.12)
@@
    Now we prove $\text{IH}_{\tilde{C}^{\tilde{N}-\tilde{T}}}-$(i) from time $T+n_{k}^{(j)}+N+1$ up to time $T+n_{k}^{(j)}+N+n_{0,k}^{(j)}$. Indeed, by $\text{IH}_{Q}$, (8.0.12) and (8.0.10), for all $1\leq s\leq n_{0,k}^{(j)}$, we get,
    $\text{IH}_{\tilde{C}}\text{-(i)P5: }\left|Df^{T+n_{k}^{(j)}+N+s}_{x}v\right|$ $\geq$ $\mu^{s}v^{(4)}_{2}\geq\mu^{s}\frac{1}{4K^{3}}\left(\lambda^{\theta}\mu\right)^{n_{k}^{(j)}}\left|Df^{T+n_{k}^{(j)}+N}_{x}v\right|$
    $\geq$ $\frac{C}{4K^{5}}\cos\varphi\cdot\mu^{s+n_{k}^{(j)}}\rho^{T}\left(\lambda^{\theta}\mu\right)^{n_{k}^{(j)}}\left|v\right|$
    $\geq$ $C\rho^{T+n_{k}^{(j)}+N+s}\left|v\right|.$
@@
    $\frac{1}{4K^{5}}\cos\varphi\left(\frac{\mu}{\rho}\right)^{n_{k}^{(j)}+s}\left(\lambda^{\theta}\mu\right)^{n_{k}^{(j)}}\geq\rho^{N},\text{ }\forall 1\leq s\leq n_{0,k}^{(j)}$
@@
    $\frac{1}{4K^{5}}\cos\varphi\left(\frac{\mu}{\rho}\right)\left(\lambda^{\theta}\mu^{2}\right)^{n_{k}^{(j)}}\geq\rho^{n_{k}^{(j)}+N}.$
    This follow from the condition imposed in (8)-(oo) and taking $n_{k}^{(j)}$ large enough. In particular, $\text{IH}_{\tilde{C}}-$(i)P5 proves
    $\left|Df^{T+n_{k}^{(j)}+N+n_{0,k}^{(j)}}_{x}v\right|\geq\frac{C}{4K^{5}}\cos\varphi\cdot\mu^{n_{0,k}^{(j)}}\rho^{T}\left(\lambda^{\theta}\mu^{2}\right)^{n_{k}^{(j)}}\left|v\right|.$ (8.0.13)
    Moreover, by (8.0.11),
    $\tan\angle\left(Df_{x}^{T+n_{k}^{(j)}+N}v,e_{2}\right)\leq\frac{2K^{3}}{\left(\lambda^{\theta}\mu\right)^{n_{k}^{(j)}}},$
    which together with (8.0.2) and by taking $n_{k}^{(j)}$ large enough imply that,
    $\tan\angle\left(Df_{x}^{T+n_{k}^{(j)}+N+n_{0,k}^{(j)}}v,e_{2}\right)\leq\left(\frac{\lambda}{\mu}\right)^{n_{0,k}^{(j)}}\frac{2K^{3}}{\left(\lambda^{\theta}\mu\right)^{n_{k}^{(j)}}}<<1.$ (8.0.14)
@@
    We start proving $\text{IH}_{\tilde{C}^{\tilde{N}-\tilde{T}}}-$(i) from time $T+n_{k}^{(j)}+N+n_{0,k}^{(j)}+1$ up to time $T+n_{k}^{(j)}+N+n_{0,k}^{(j)}+M$. Indeed, by $\text{IH}_{H}$ and (8.0.13), for all $1\leq s\leq M$, we get,
    $\text{IH}_{\tilde{C}}\text{-(i)P6: }\left|Df^{T+n_{k}^{(j)}+N+n_{0,k}^{(j)}+s}_{x}v\right|$ $\geq$ $\frac{1}{K}\left|Df^{T+n_{k}^{(j)}+N+n_{0,k}^{(j)}}_{x}v\right|$
    $\geq$ $\frac{C}{4K^{6}}\cos\varphi\cdot\mu^{n_{0,k}^{(j)}}\rho^{T}\left(\lambda^{\theta}\mu^{2}\right)^{n_{k}^{(j)}}\left|v\right|$
    $\geq$ $C\rho^{T+n_{k}^{(j)}+N+n_{0,k}^{(j)}+s}\left|v\right|.$